OpenStudy (anonymous):

Solve 2x + 5y = –13 3x – 4y = –8 a. (4, 1) b. (–4, 1) c.(4, –1) d.(–4, –1)

OpenStudy (anonymous):

do you want the answer or how to solve it ? :P

OpenStudy (inkyvoyd):

There's an easy way and a hard way to do this problem.

OpenStudy (inkyvoyd):

The first way is to substitute like an idiot. Works, but can be slower. (if one ordered pair matches both equations then that is your answer)

OpenStudy (anonymous):

no answe i want to learn how to solve my self please

OpenStudy (inkyvoyd):

Second way is to multiply equation one by 3, and equation two by 2, and subtract the equations.

OpenStudy (anonymous):

@inkyvoyd why do you call idiots the one who substitute?? don't be rude dude!

sam (.sam.):

2x + 5y = –13 3x – 4y = –8 From (1) 2x + 5y = –13 x = (–13-5y)/2 --------------------- 3x – 4y = –8 3 (–13-5y)/2 – 4y = –8 (3(-5y-13))/(2)-4y=-8 (3(-5y-13))/(2)-4y*(2)/(2)=-8 (3(-5y-13))/(2)-(8y)/(2)=-8 (3(-5y-13)-(8y))/(2)=-8 (-23y-39)/(2)=-8 -23y-39=-16 -23y=23 y=-1 --------------------------------- x = (–13-5y)/2 x = (–13-5(-1))/2 x=(-13+5)/(2) x=(-8)/(2) x=-4 (-4,-1)

OpenStudy (inkyvoyd):

Well, sam wins.

OpenStudy (anonymous):

I still want to learn how to solve onn my own no answers

OpenStudy (inkyvoyd):

And, idiots are ones who subsitute because if the coefficients of the equation are very nasty, they will have much trouble solving. Any big numbers, or string of radicals will make it very hard to try reverse substitution to solve.

OpenStudy (inkyvoyd):

The ones where substitution are fast I can usually solve in my head with trial and error.

OpenStudy (inkyvoyd):

btw, there's a third way, called actual "subsitution"

OpenStudy (inkyvoyd):

But sam posted that too, nvm

OpenStudy (anonymous):

could you explain like the first step into solve because i always have a hard time with this type of problems